The perimeter of a rectangle is 52m. The length is 8 more than twice the width. Find the dimensions...what is the length and what is the width?
Answer by edjones(7311)

L=2w+8
P=2L+2w
2(2w+8)+2w=52
4w+16+2w=52
6w+16=52
6w=36
w=6m
L=2*6+8=20m
.
Ed